OpenStudy (anonymous):

last question for today...it's about application in economics.. suppose a company has a marginal cost function y = C' x√(9+x²), where x is the number of thousands of items sold and the cost C is in thousand of dollar. if fixed cost are $11 thousand, what is the total cost of manufacturing the first 4 thousand items? I used intergration to find the C(x) then I substitute x=4 and C= 11.. is that correct?

OpenStudy (anonymous):

C'= x√(9+x²) you mean?

OpenStudy (anonymous):

that is marginal cost

OpenStudy (anonymous):

sorry.. c'(x)=x√(9+x²) yeah..

OpenStudy (anonymous):

yes, integrate marginal cost to get total cost; use u-sub

OpenStudy (anonymous):

is that correct when i subs C=11?

OpenStudy (anonymous):

integrated out to be 1/3 (x^2+9)^(3/2)+c

OpenStudy (anonymous):

1/3 (4^2+9)^(3/2)+11

OpenStudy (anonymous):

i got 2/3 (9 + x^2)^(3/2) + C

OpenStudy (anonymous):

u=x^2+9 du=2x dx \[1/2\int \sqrt{u} du\]

OpenStudy (anonymous):

oh.. i forgot that.. thanks.. (: